• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ramesh

    Log entry time 23:57:06 on May17,2008

    Entry number 238530

    keyword=Shift summary

    16:00- 18:20 Folks in the back room are troublshooting the reason for high deadtime (33%) even at 3 uA beam by frequently killing and reviving the CODA. Also trying to understand whether the BigBite position (-32 deg) was good enough to catch the elastic events without clipping (this clipping of elastic events on the BigBite has been an issue for sometime).

    17:15 - 17:30 Target high power heater was shut off by itself, Paul Reimer (TO) brought it back by changing the PID mode to first a manual and next (once it was working ok) the automatic mode.

    18:20 - 19:15 Started a new test kinematics (left hrs angle 56.936 deg, momentum setting = 0.6318 MeV/c) setting to check whether we were clipping the elastic events on BigBite.

    19:15- 20:00 Started another test kinematics (left hrs angle 54.4 deg, momentum setting = 0.6837 MeV/c) setting to again check whether we were clipping the elastic events on BigBite.

    19:30 Moved the target to empty position.

    20:00 Started the production kin (del = 0: left hrs angle = 57.5 deg, momentum setting = 0.6162 GeV/c) finally!

    21:05 - 21:25 MCC reported having RF problem, beam stops.

    Continueing production data on LH2 with rastor ON even at 3uA.


    22:47 - 24:00 Asked for escorted access to take a look into the BigBite electronics. Still on access while the shift ends.